# Rethinking Circuit Completeness in Language Models: AND, OR, and ADDER Gates

### v1

@@ 1 Introduction
As an intervention-based approach to mechanistic interpretability, circuit discovery allows for the extraction of subgraphs from the computational graph of a language model that play a significant role in task performance, referred to as circuits (Elhage et al., 2021; Conmy et al., 2023; Rai et al., 2024; Olah et al., 2020). Several key studies have supported its development (Hsu et al., 2024; Haklay et al., 2025), such as those focusing on ensuring that circuits faithfully reflect the model’s outputs (Conmy et al., 2023; Marks et al., 2024), enabling efficient circuit extraction (Syed et al., 2024), and addressing scalability challenges for models with extremely large parameters and corpora (Yu et al., 2024; Bhaskar et al., 2024; Lieberum et al., 2023).
As the concept of circuits evolves, recent attention has increasingly focused on the completeness of circuits in addition to faithfulness. For example, completeness has been redefined such that when a circuit is removed from the computational graph, the performance of the task should degrade significantly (De Cao et al., 2022; Bayazit et al., 2023). Nevertheless, current circuit discovery methods have been found to lack completeness (Yu et al., 2024). Moreover, theory analysis (Mueller, ) indicates that incomplete circuits lead to two potential pitfalls: non-transitivity and preemption, which prevent the recovery of the key mechanisms underlying the circuit. Finally, incompleteness results in variability in circuit discovery outcomes, making the circuit appear more like an arithmetic solution to obtain the output rather than representing a closed-form solution with interpretability (Chen et al., 2024).
Incompleteness largely arises from the presence of OR gates (Wang et al., ; Conmy et al., 2023). For instance, consider a model $M$ that employs two identical and disjoint serial circuit paths, $C_{1}$ and $C_{2}$, which operate in parallel and whose outputs are subsequently combined via an OR operation. In this case, identifying either path is sufficient to achieve faithfulness, and removing the other path is a preferable choice for promoting sparsity. However, restoring completeness by discovering OR gates remains a challenge. The simplest approach, which involves repeated interventions on combinations of components (Mueller, ), could theoretically uncover the complete OR gate; however, this makes circuit discovery an NP problem. Additionally, while denoising-based intervention methods can rapidly restore OR gates (Heimersheim and Nanda, 2024), they lead to a more severe loss of faithfulness. Furthermore, these methods fail to isolate the OR gates from the final circuit, resulting in a lack of logical interpretability.
@@
1. We systematically introduce three types of logic gates that compose a circuit: AND, OR, and ADDER gates. Through these gates, we are able to infer the minimum requirements for a circuit to achieve faithfulness and completeness, as well as assess the capability of noising-based and denoising-based interventions in restoring these gates. Based on these corollaries, we analyze three types of prevailing circuit discovery methods, named greedy search (Conmy et al., 2023; Yao et al., ; Lieberum et al., 2023), linear estimation (Syed et al., 2024; Nanda, 2023), and differentiable mask (Yu et al., 2024; De Cao et al., 2022; Bhaskar et al., 2024), by evaluating their ability to recover the three logic gates and their faithfulness and completeness. Moreover, we conduct experiments to provide empirical evidence supporting these theoretical conclusions.
@@ 2.1 Circuit Discovery
In Transformer decoder-based language models, the forward pass is typically conceptualized as a computational graph $\mathcal{G}$, where the nodes represent components (such as attention heads, MLPs, or even more granular elements like the query, key, and value matrices) and an edge $i\rightarrow j$ denotes a connection where the output of component $i$ serves as input to component $j$. Circuit discovery seeks to identify a subgraph (circuit) $\mathcal{C}\subset\mathcal{G}$ that captures the task-relevant behavior of the model (Rai et al., 2024).
The process used to prune and obtain the circuit $\mathcal{C}$ is referred to as intervention (also known as knockout, ablation(Heimersheim and Nanda, 2024; Vig et al., 2020; Chan et al., 2022; Goldowsky-Dill et al., 2023). For a given task $\mathcal{T}$, each sample $x$ is referred to as clean text, and the corresponding forward pass yields the clean activation $x_{i}$ at each component $i$. A perturbed version of the input, denoted $\tilde{x}$, is called corrupted text, producing a corresponding corrupted activation (Zhang and Nanda, 2023; Heimersheim and Nanda, 2024). The corrupted activation $\tilde{x}_{i}$ depends on the specific ablation method used. For example, zero ablation sets $\tilde{x}_{i}=0$, while noise ablation draws $\tilde{x}_{i}$ from a predefined noise distribution. A widely used method, interchange ablation, defines $\tilde{x}_{i}$ as the activation resulting from an input text that has been minimally perturbed to produce a different task label (Bhaskar et al., 2024).
The intervention is divided into two strategies: noising-based intervention (hereafter referred to as Ns) and denoising-based intervention (hereafter referred to as Dn(Meng et al., 2022). The Ns first runs the clean text in the computational graph. Then, corrupted activations replace each clean activation to observe the change in the final output $y$. If replaced (also known as removed or pruned) activations lead to a significant change in output, they are considered to make an important contribution to the task $\mathcal{T}$ and should be retained in the circuit $\mathcal{C}$ (Heimersheim and Nanda, 2024). Let $p_{\mathcal{G}}(y|x)$ denote the model’s original output, $p_{\mathcal{C}}(y|x,\tilde{x})$ represent the circuit’s output after intervention. Specifically, if an edge $j\to i$ is retained within $\mathcal{C}$, the activation of component $i$ keeps the clean one ($x_{i}$). Conversely, it is replaced by the corrupted one ($\tilde{x}_{i}$). Let $s$ denote the requirement of sparsity, and $D$ represent the distance used to quantify the difference between the two outputs. Ns has the following objective:
@@
Most of the related work on circuit discovery follows the Ns strategy. We categorize these works into three types: (1) Greedy search (Conmy et al., 2023; Yao et al., ; Lieberum et al., 2023), which iteratively examines each edge (or node) through intervention to obtain a greedy solution for the circuit. (2) Linear estimation (Syed et al., 2024; Nanda, 2023), where the contribution of each edge is approximated by a gradient measure obtainable in a single backward pass. This approach ranks the importance of each edge to approximate the circuit. (3) Differentiable masks (Yu et al., 2024; De Cao et al., 2022; Bhaskar et al., 2024), where a learnable mask is assigned to each edge (or node), treating circuit discovery as an optimization problem to derive the optimal circuit.
@@ 2.2 Circuit Evaluation
Faithfulness refers to the circuit’s ability to perform task $\mathcal{T}$ in isolation, which is defined as the difference between the circuit’s output and the model’s original output (Wang et al., ; Yu et al., 2024; Heimersheim and Nanda, 2024). This is represented in Equations 1 as $\mathbb{E}_{(x,\tilde{x})\in\mathcal{T}}[D(p_{\mathcal{G}}(y|x)||p_{\mathcal{C}}(y|x,\tilde{x}))]$ (simplified as $D(\mathcal{G}||\mathcal{C})$). Method ACDC (Wang et al., ) measures faithfulness by computing the average difference in the unnormalized output logits between the correct token and an incorrect option. Recently, work (Conmy et al., 2023; Heimersheim and Nanda, 2024; Kim et al., 2021) proposes that KL divergence provides a better measure of the distribution over the vocabulary, while other work (Yu et al., 2024; Chen et al., 2024) suggests that task accuracy can avoid the overemphasis on irrelevant vocabulary in the KL divergence. In this paper, we measure faithfulness using both KL divergence and task accuracy as metrics.
Completeness refers to whether the circuit includes all the important paths that have an effect on the output. The work (Wang et al., ) first introduces the concept of circuit completeness, stating that $\mathcal{C}$ and $\mathcal{G}$ should ensure similar outputs even under any knockout. Therefore, the incompleteness score is defined as the difference $D(\mathcal{C}\setminus\mathcal{K}||\mathcal{G}\setminus\mathcal{K})$ for any subcircuit $\mathcal{K}\subset\mathcal{C}$. Existing work (Yu et al., 2024; Chen et al., 2024) proposes that insufficient sampling of $\mathcal{K}$ may lead to unreliable approximations, and thus recommends evaluating completeness by assessing the performance after the circuit’s removal from the computational graph on the task $\mathcal{T}$, i.e., $D(\mathcal{G}\setminus\mathcal{C}||\mathcal{G})$ (Bayazit et al., 2023; De Cao et al., 2022). In this paper, we also adopt it to evaluate completeness.
Sparsity refers to that the circuit should be as small as possible. Currently, many works (Bhaskar et al., 2024; Yu et al., 2024; Chen et al., 2024) recommend measuring sparsity using the ratio $|\mathcal{C}|/|\mathcal{G}|$, which represents the proportion of edges in the circuit relative to those in the computational graph. In fact, higher sparsity tends to result in lower faithfulness, meaning that the circuit always reflects some trade-off between sparsity and faithfulness.
@@ 3.1 Logical Gates
To better analyze the faithfulness and completeness of circuits, we systematically introduce three fundamental circuit logic types: the AND gate, OR gate, and ADDER gate.
@@
For any edge $i\rightarrow j$, node $j$ is referred to as the receiver node, and node $i$ is referred to as the sender node. The logically complete circuit usually contains:
AND gate: There exists a receiver node $B$, which is connected by more than 1 sender node $A_{1},A_{2},\dots$, and all sender nodes satisfy an AND logical relationship with the receiver node, i.e., $B=A_{1}\land A_{2}\land\dots$. In this case, the set $\{(A_{1},A_{2},\dots),B\}$ forms an AND gate.
@@
OR gate: There exists a receiver node $B$, which is connected by more than one sender node $A_{1},A_{2},\dots$, and all sender nodes satisfy an OR logical relationship with the receiver node, i.e., $B=A_{1}\lor A_{2}\lor\dots$. In this case, the set $\{(A_{1},A_{2},\dots),B\}$ forms an OR gate.
ADDER gate: There exists a receiver node $B$, which is connected by one or more sender nodes $A_{1},\dots$, and all sender nodes satisfy an ADDER logical relationship with the receiver node, i.e., $B=A_{1}+A_{2}+\dots$. In this case, the set $\{(A_{1},\dots),B\}$ forms an ADDER gate.
For example, let the set $\{(A_{1},A_{2}),B\}$ be the toy circuit, and $B$ is connected to the output [^1] If there are additional gates between the receiver node and the output, it becomes difficult to draw isolated conclusions and understand the effects of each gate. Therefore, for analytical convenience, all analyses in this section assume that the receiver node is directly connected to the output.. If the set $\{(A_{1},A_{2}),B\}$ forms an AND gate, i.e., $B=A_{1}\land A_{2}$, then $B$ can influence the result only when both $A_{1}$ and $A_{2}$ are present in the circuit. If either $A_{1}$ or $A_{2}$ (or both) are removed, $B$ will no longer affect the result [^2] In practice, $B=A_{1}$ may result in a negligible effect, which can be considered insignificant. Similarly, when an edge is removed from an OR gate, it may also lead to a change that is so small as to be ignored.. Similarly, if the set $\{(A_{1},A_{2}),B\}$ forms an OR gate, i.e., $B=A_{1}\lor A_{2}$, then $B$ can consistently influence the result when either $A_{1}$ or $A_{2}$ (or both) are present in the circuit. Only when both $A_{1}$ and $A_{2}$ are removed can $B$ cease to affect the result. If the set $\{(A_{1},A_{2}),B\}$ forms an ADDER gate, i.e., $B=A_{1}+A_{2}$, then both $A_{1}$ and $A_{2}$ contribute significantly to $B$ in an additive manner. When either $A_{1}$ or $A_{2}$ is removed, the effect of $B$ on the result decreases in isolation. Moreover, we design a toy model to study ADD, OR, ADDER gates in Appendix C.
@@
By Definition 1, any circuit can be represented as a combination of AND, OR, and ADDER gates. We can draw a corollary regarding Noising-based Intervention (Ns) and Denoising-based Intervention (Dn) through the use of AND, OR, and ADDER gates:
@@
Ns is responsible for recovering the complete AND and ADDER gates, but cannot recover the complete OR gates. Dn is responsible for recovering the complete OR and ADDER gates, but cannot recover the complete AND gates (The proofs are shown in Appendix A).
Corollary 1 demonstrates the performance of both Ns and Dn on different logical gates (some of the conclusions are also supported in the work (Heimersheim and Nanda, 2024)). Given that current circuit discovery methods predominantly rely on Ns, these methods are unable to fully recover the OR gate. A detailed analysis of this limitation is provided in Section 3.2.
Additionally, these logical gates reveal some interesting phenomena as shown in Figure 1. For optimal faithfulness and sparsity, the circuit only needs to include one edge from each OR gate. For optimal completeness and sparsity, the circuit only needs to include one edge from each AND gate. Based on the definitions in Section 2, we can draw the following corollary regarding these properties:
Corollary 2.
The minimal edge subset that satisfies optimal faithfulness consists of all edges from the AND gates, all edges from the ADDER gates, and any one edge from each OR gate. The minimal edge subset that satisfies optimal completeness consists of all edges from the OR gates, all edges from the ADDER gates, and any one edge from each AND gate (The proofs are shown in Appendix B).
Corollary 2 provides insights for better understanding these three properties. In any gate, the influence of the receiver node on the output can be regarded as the “gate effect” (The collective gate effects ensure that the circuit $\mathcal{C}$ approximates the functionality of computation graph $\mathcal{G}$.) On the top of this, faithfulness refers to the sum of all gate effects in the circuit, which should be as large as possible (i.e., the closer it is to the functionality of $\mathcal{G}$, the better); completeness refers to that when the circuit is removed, the sum of all gate effects should be as small as possible (i.e., the greater the deviation from the functionality of $\mathcal{G}$, the better); sparsity refers to that, while striving to maximize both faithfulness and completeness, the number of edges in the circuit should be as minimal as possible.
3.2 Logical Analysis of Circuit Discovery
Based on Corollary 1, we can combine the specific types of circuit discovery to determine their capabilities across the three types of logic gates. Building on this, according to Corollary 2, we can further evaluate their performance in terms of faithfulness and completeness. Table 1 presents the specific results for the three types of circuit discovery methods mentioned in Section 2.
@@ 3.2 Logical Analysis of Circuit Discovery
linear estimation (Syed et al., 2024; Nanda, 2023) $\surd$ $\times$ $\surd$ $\times$ $\times$
@@
linear estimation (Syed et al., 2024; Nanda, 2023) $\times$ $\surd$ $\surd$ $\times$ $\times$
@@
According to Corollary 1, methods from Ns are able to identify complete AND and ADDER gates. Among these, methods based on greedy search and differentiable masks can identify partial OR gates, whereas methods based on linear estimation are unable to detect any edges of OR gates. Similarly, methods from Dn exhibit a similar pattern. While they can completely identify OR and ADDER gates, methods based on greedy search and differentiable masks can detect partial AND gates, while methods based on linear estimation fail to identify any. In Appendix C, we explain why greedy search and differentiable mask methods are able to identify some edges, whereas linear estimation completely fails to do so. Moreover, inspired by (Conmy et al., 2023), we design a simple one-layer transformer toy model to implement the basic AND, OR, and ADDER gates, and validate the performance of these circuit discovery methods corresponding to the conclusion from Table 1.
4 Complete Discovery of Logical Gates
4.1 Separating AND, OR, and ADDER Gates
We denote the circuit constructed under the Ns strategy as $\mathcal{C}_{\text{Ns}}$, and the one constructed under the Dn strategy as $\mathcal{C}_{\text{Dn}}$. Based on the set-theoretic relationships between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$ (see Corollary 1), we extract subsets of edges corresponding to AND, OR, and ADDER gates as follows:
We conduct an ablation on these edges: for each gate, we randomly remove either one or two edges on the same receiver node and measure the resulting change in the KL divergence of the output. This procedure is repeated 30 times for each receiver node, and the distributions of $\Delta\mathrm{KL}$ values are summarized via box plots, as shown in Figure 3. We selected the computational graph of GPT2-small as $\mathcal{G}$, and Indirect Object Inference (IOI) (Wang et al., ) as the test task. For the baseline methods, we chose ACDC (Conmy et al., 2023) to represent the greedy search method, EAP (Syed et al., 2024) to represent the linear estimation method, and EdgePruning (Bhaskar et al., 2024) to represent the differentiable mask method. For details regarding the implementation of these strategies within each baseline, we refer the reader to Appendix D.

### v2

@@ 1 Introduction
As an intervention-based approach to mechanistic interpretability, circuit discovery allows for the extraction of subgraphs from the computational graph of a language model that play a significant role in task performance, referred to as circuits (elhage2021mathematical, ; conmy2023towards, ; rai2024practical, ; olah2020zoom, ). Several key studies have supported its development (hsu2024efficient, ; haklay2025position, ), such as those focusing on ensuring that circuits faithfully reflect the model’s outputs (conmy2023towards, ; marks2024sparse, ), enabling efficient circuit extraction (syed2024attribution, ), and addressing scalability challenges for models with extremely large parameters and corpora (yu2024functional, ; bhaskar2024finding, ; lieberum2023does, ).
As the concept of circuits evolves, recent attention has increasingly focused on the completeness of circuits in addition to faithfulness. For example, completeness has been redefined such that when a circuit is removed from the computational graph, the performance of the task should degrade significantly (de2022sparse, ; bayazit2023discovering, ). Nevertheless, current circuit discovery methods have been found to lack completeness (yu2024functional, ). Moreover, theory analysis (mueller2024missed, ) indicates that incomplete circuits lead to two potential pitfalls: non-transitivity and preemption, which prevent the recovery of the key mechanisms underlying the circuit. Finally, incompleteness results in variability in circuit discovery outcomes, making the circuit appear more like an arithmetic solution to obtain the output rather than representing a closed-form solution with interpretability (chen2024unveiling, ).
Incompleteness largely arises from the presence of OR gates (wanginterpretability, ; conmy2023towards, ). For instance, consider a model $M$ that employs two identical and disjoint serial circuit paths, $C_{1}$ and $C_{2}$, which operate in parallel and whose outputs are subsequently combined via an OR operation. In this case, identifying either path is sufficient to achieve faithfulness, and removing the other path is a preferable choice for promoting sparsity. However, restoring completeness by discovering OR gates remains a challenge. The simplest approach, which involves repeated interventions on combinations of components (mueller2024missed, ), could theoretically uncover the complete OR gate; however, this makes circuit discovery an NP problem (We explained it in Appendix E.1). Additionally, while denoising-based intervention methods can rapidly restore OR gates (heimersheim2024use, ), they lead to a more severe loss of faithfulness. Furthermore, these methods fail to isolate the OR gates from the final circuit, resulting in a lack of logical interpretability.
@@
1. We systematically introduce three types of logic gates that compose a circuit: AND, OR, and ADDER gates. Through these gates, we are able to infer the minimum requirements for a circuit to achieve faithfulness and completeness, as well as assess the capability of noising-based and denoising-based interventions in restoring these gates. Based on these corollaries, we analyze three types of prevailing circuit discovery methods, named greedy search (conmy2023towards, ; yaoknowledge, ; lieberum2023does, ), linear estimation (syed2024attribution, ; nanda2023attribution, ), and differentiable mask (yu2024functional, ; de2022sparse, ; bhaskar2024finding, ), by evaluating their ability to recover the three logic gates and their faithfulness and completeness. Moreover, we conduct experiments to provide empirical evidence supporting these theoretical conclusions.
@@ 2.1 Circuit Discovery
In Transformer decoder-based language models, the forward pass is typically conceptualized as a computational graph $\mathcal{G}$, where the nodes represent components (such as attention heads, MLPs, or even more granular elements like the query, key, and value matrices) and an edge $i\rightarrow j$ denotes a connection where the output of component $i$ serves as input to component $j$. Circuit discovery seeks to identify a subgraph (circuit) $\mathcal{C}\subset\mathcal{G}$ that captures the task-relevant behavior of the model (rai2024practical, ).
The process used to prune and obtain the circuit $\mathcal{C}$ is referred to as intervention (also known as knockout, ablation(heimersheim2024use, ; vig2020investigating, ; chan2022causal, ; goldowsky2023localizing, ). For a given task $\mathcal{T}$, each sample $x$ is referred to as clean text, and the corresponding forward pass yields the clean activation $x_{i}$ at each component $i$. A perturbed version of the input, denoted $\tilde{x}$, is called corrupted text, producing a corresponding corrupted activation (zhang2023towards, ; heimersheim2024use, ). The corrupted activation $\tilde{x}_{i}$ depends on the specific ablation method used. For example, zero ablation sets $\tilde{x}_{i}=0$, while noise ablation draws $\tilde{x}_{i}$ from a predefined noise distribution. A widely used method, interchange ablation, defines $\tilde{x}_{i}$ as the activation resulting from an input text that has been minimally perturbed to produce a different task label (bhaskar2024finding, ).
[…]
@@
Most of the related work on circuit discovery follows the Ns strategy. We categorize these works into three types: (1) Greedy search (conmy2023towards, ; yaoknowledge, ; lieberum2023does, ), which iteratively examines each edge (or node) through intervention to obtain a greedy solution for the circuit. (2) Linear estimation (syed2024attribution, ; nanda2023attribution, ), where the contribution of each edge is approximated by a gradient measure obtainable in a single backward pass. This approach ranks the importance of each edge to approximate the circuit. (3) Differentiable masks (yu2024functional, ; de2022sparse, ; bhaskar2024finding, ), where a learnable mask is assigned to each edge (or node), treating circuit discovery as an optimization problem to derive the optimal circuit.
@@ 2.2 Circuit Evaluation
Faithfulness refers to the circuit’s ability to perform task $\mathcal{T}$ in isolation, which is defined as the difference between the circuit’s output and the model’s original output (wanginterpretability, ; yu2024functional, ; heimersheim2024use, ). This is represented in Equations 1 as $)]\mathbb{E}_{(x,\tilde{x})\in\mathcal{T}}[D(p_{\mathcal{G}}(y|x)||p_{\mathcal{C}}(y|x,\tilde{x}$))] (simplified as $D(\mathcal{G}||\mathcal{C})$). Method ACDC (wanginterpretability, ) measures faithfulness by computing the average difference in the unnormalized output logits between the correct token and an incorrect option. Recently, work (conmy2023towards, ; heimersheim2024use, ; kim2021comparing, ) proposes that KL divergence provides a better measure of the distribution over the vocabulary, while other work (yu2024functional, ; chen2024unveiling, ) suggests that task accuracy can avoid the overemphasis on irrelevant vocabulary in the KL divergence. In this paper, we measure faithfulness using both KL divergence and task accuracy as metrics.
Completeness refers to whether the circuit includes all the important paths that have an effect on the output. The work (wanginterpretability, ) first introduces the concept of circuit completeness, stating that $\mathcal{C}$ and $\mathcal{G}$ should ensure similar outputs even under any knockout. Therefore, the incompleteness score is defined as the difference $D(\mathcal{C}\setminus\mathcal{K}||\mathcal{G}\setminus\mathcal{K})$ for any subcircuit $\mathcal{K}\subset\mathcal{C}$. Existing work (yu2024functional, ; chen2024unveiling, ) proposes that insufficient sampling of $\mathcal{K}$ may lead to unreliable approximations (we show the practical results in Appendix A), and thus recommends evaluating completeness by assessing the performance after the circuit’s removal from the computational graph on the task $\mathcal{T}$, i.e., $D(\mathcal{G}\setminus\mathcal{C}||\mathcal{G})$ (bayazit2023discovering, ; de2022sparse, ). In this paper, we also adopt it to evaluate completeness.
Sparsity refers to that the circuit should be as small as possible. Currently, many works (bhaskar2024finding, ; yu2024functional, ; chen2024unveiling, ) recommend measuring sparsity using the ratio $|\mathcal{C}|/|\mathcal{G}|$, which represents the proportion of edges in the circuit relative to those in the computational graph. In fact, higher sparsity tends to result in lower faithfulness, meaning that the circuit always reflects some trade-off between sparsity and faithfulness.
@@ 3.1 Logical Gates
Recent studies (wanginterpretability, ; conmy2023towards, ; heimersheim2024use, ) have increasingly observed that within a circuit, certain subcircuits influence the output according to logical relationships resembling AND or even OR operations. Building on these findings, we systematically introduce three fundamental circuit logic types: the AND gate, OR gate, and ADDER gate.
@@
We assume a common paradigm in which a receiver node $B$, which is connected by more than 1 sender node $A_{1},A_{2},$.... For any edge $A_{i}\rightarrow B$, we use binary values ‘0’ and ‘1’ to represent the activation state of a node. Specifically, $A_{i}=0$ indicates that node $A_{i}$ is removed, ablated, or deactivated, whereas $A_{i}=1$ indicates that node $A_{i}$ is retained and active. When the sender nodes are ablated, the effect of node $B$ on the output exhibits three distinct patterns, which are as follows:
AND: All sender nodes satisfy an AND logical relationship with the receiver node, i.e., $B=A_{1}\land A_{2}\land\dots$. In this case, node $B$ exerts a significant effect on the output only if all of its sender nodes are retained. If even a single sender node is ablated, the effect of $B$ on the output is nearly eliminated.
@@
OR gate: All sender nodes satisfy an OR logical relationship with the receiver node, i.e., $B=A_{1}\lor A_{2}\lor\dots$. In this case, node $B$ always exerts a significant effect on the output if one or more of its sender nodes are retained. Only if all sender nodes are ablated, the effect of $B$ on the output is nearly eliminated.
ADDER gate: all sender nodes satisfy an ADDER logical relationship with the receiver node, i.e., $B=A_{1}+A_{2}+\dots$. In this case, node $B$ exhibits its maximal effect on the output only when all of its sender nodes are retained. If any single sender node is ablated, the effect of $B$ on the output is substantially diminished; when all sender nodes are ablated, $B$’s effect on the output is reduced to zero. Accordingly, we define the state of $B$ as taking values 0,1,2,…, where the total number of distinct states equals the number of sender nodes.
Theoretical analyses support the view that Ns is capable of recovering a complete AND gate but fails to recover a complete OR gate, whereas Dn demonstrates the opposite pattern (heimersheim2024use, ). This asymmetry is straightforward to interpret. The Ns procedure corresponds to the transition from a clean activation state ($\text{state}=1$) to a corrupted activation state ($\text{state}=0$). Since all gates can be regarded as being initialized with activation states equal to $1$, any transition to $\text{state}=0$ induces a significant change in the effect of AND and ADDER gates on the output. Consequently, Ns can reliably identify AND and ADDER gates. Conversely, the Dn procedure corresponds to initialization with activation states equal to 0. In this case, any transition to $\text{state}=1$ produces a significant change in the effect of OR and ADDER gates on the output. Moreover, we design a toy model to explain ADD, OR, ADDER gates in Appendix D.
Therefore, we denote the circuit constructed under the Ns strategy as $\mathcal{C}_{\text{Ns}}$, and the one constructed under the Dn strategy as $\mathcal{C}_{\text{Dn}}$. Based on the above set-theoretic relationships between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$, we extract subsets of edges corresponding to AND, OR, and ADDER gates as follows:
We conduct an ablation on these edges: for each gate, we randomly remove either one or two edges on the same receiver node and measure the resulting change in the KL divergence of the output. This procedure is repeated 30 times for each receiver node, and the distributions of $\Delta\mathrm{KL}$ values are summarized via box plots, as shown in Figure 1 (Detailed results are shown in Appendix E.2). We selected the computational graph of GPT2-small as $\mathcal{G}$, and Indirect Object Inference (IOI) (wanginterpretability, ) as the test task. For the baseline methods, we chose ACDC (conmy2023towards, ) to represent the greedy search method, EAP (syed2024attribution, ) to represent the linear estimation method, and EdgePruning (bhaskar2024finding, ) to represent the differentiable mask method. For details regarding the implementation of these strategies within each baseline, we refer the reader to Appendix E.
Figure 1: $\Delta KL$ in removing 1 and 2 edges of AND, OR, ADDER gates.
@@
3.2 How to Use Logical Gates to Interpret Faithfulness and Completeness
These logical gates reveal some interesting phenomena as shown in Figure 2. For optimal faithfulness and sparsity, the circuit only needs to include one edge from each OR gate. For optimal completeness and sparsity, the circuit only needs to include one edge from each AND gate. Based on the definitions in Section 2, we can draw the following corollary regarding these properties:
@@ 3.2 How to Use Logical Gates to Interpret Faithfulness and Completeness
The minimal edge subset that satisfies optimal faithfulness consists of all edges from the AND gates, all edges from the ADDER gates, and any one edge from each OR gate. The minimal edge subset that satisfies optimal completeness consists of all edges from the OR gates, all edges from the ADDER gates, and any one edge from each AND gate (The proofs are shown in Appendix C).
Corollary 1 provides an explanation for why existing circuit discovery algorithms predominantly adopt the Ns strategy. The reason is that Ns is able to recover complete AND and ADDER gates, while its recovery of OR gates remains incomplete. This tradeoff corresponds precisely to the optimal balance between faithfulness and sparsity. Conversely, the Dn strategy is not adopted because it fails to recover complete AND gates, thereby severely compromising the faithfulness of the resulting circuit.
Therefore, we further evaluate the performance of existing work in terms of faithfulness and completeness. Table 1 presents the specific results for the three types of circuit discovery methods mentioned in Section 2.
@@
linear estimation (syed2024attribution, ; nanda2023attribution, ) $\surd$ $\times$ $\surd$ $\times$ $\times$
@@
linear estimation (syed2024attribution, ; nanda2023attribution, ) $\times$ $\surd$ $\surd$ $\times$ $\times$
@@
Among these, methods based on greedy search and differentiable masks can identify partial OR gates, whereas methods based on linear estimation are unable to detect any edges of OR gates. Similarly, methods from Dn exhibit a similar pattern. While they can completely identify OR and ADDER gates, methods based on greedy search and differentiable masks can detect partial AND gates, while methods based on linear estimation fail to identify any. In Appendix D, we explain why greedy search and differentiable mask methods are able to identify some edges, whereas linear estimation completely fails to do so. Moreover, inspired by (conmy2023towards, ), we design a simple one-layer transformer toy model to implement the basic AND, OR, and ADDER gates, and validate the performance of these circuit discovery methods corresponding to the conclusion from Table 1.
3.3 Granularity Alignment between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$